If the distance between the earth and the sun shrinks to half the present distance then find the new duration of year.
 

Grade:10

2 Answers

Vikas TU
7 years ago
Let distance b/w the earth and the sun be x meters.
And from kepler’s law square of Time period is directly proportional to the cube of Radius of the earth.
Therefore initially,
T^2 = k*x^3........................(1)
after 
t^2 = k*(x/2)^3..............(2)
Divide them get the value of t.
Adarsh
7 years ago
Let T is the time period and r is the distance between the earth and the sun.
Then,
by Kepler’s law
    T2∝r3
=> T∝r3/2 
=> 365days ∝ r3/2 ..............(1)
When r becomes r/2
So new time period, T’ ∝ (r/2)3/2..............(2)
Eqn (2) ÷ (1) => T’ / 365 =1/ 23/2
Thus new time period T’ = 365/ 23/2 days
 
Hope you understand my solution.
 
Please approve only if my answer is correct.